OpenStudy (anonymous):

f(x)= 4x+1 and g(x)=x^2-5 find (f-g)(x) I got X^2-4x-6 and for (f+g)(x) I got x^2+4x+4 I'm a bit unsure if the subtraction is right though.

OpenStudy (anonymous):

which is being subtracted from which?

OpenStudy (anonymous):

write it out and you'll never make a mistake.

OpenStudy (anonymous):

well, almost never

OpenStudy (anonymous):

Wait I got it!

OpenStudy (mathmale):

"find (f-g)(x) I got X^2-4x-6": Be careful with your signs. First write f(x) and then g(x): f(x) = 4x + 1, g(x) = x^2 - 5 Since you are SUBTRACTING g(x) from f(x), enclose g(x) in parentheses and write a '-' sign in front of it: f(x) - g(x)

OpenStudy (mathmale):

Goood. Satisfied?

OpenStudy (anonymous):

6+4x-x^2

OpenStudy (anonymous):

I was a bit confused because of the answers none of the others started with the numerical value they all started with the x^2 I knew subtracting meant the -5 would become +6

OpenStudy (mathmale):

Looks good! (f-g)(x+ = -x^2 +4x + 6

OpenStudy (anonymous):

4x+1-x^2-(-5) makes it positive +5

OpenStudy (anonymous):

there you go but it is customary to start with the highest power of x first and then in descending power... -x^2 + 4x + 6

OpenStudy (mathmale):

...where I intentionally wrote the terms in order by descending powers of x.

OpenStudy (mathmale):

...same as @pgpilot326

OpenStudy (anonymous):

Yeah I know, thank you guys!

OpenStudy (anonymous):

lol